Slips of paper numbered 1 through 10 are placed into a bag. What is the probability of drawing the number 10 three times in a row if a slip is drawn at random and then replaced?
Answer by stanbon(75887)

Slips of paper numbered 1 through 10 are placed into a bag. What is the probability of drawing the number 10 three times in a row if a slip is drawn at random and then replaced?
------
P(draw a 10) = 1/10
-----
P(draw a 10 ten times with replacement) = (1/10)^10 = 1/10000000000
===============
Cheers,
Stan H.
===============